OpenStudy (anonymous):

Which is an everyday situation for the math expression with the context provided? ; sharing something equally

OpenStudy (anonymous):

; sharing something equally A. The Perez family drove 500 miles each day of their vacation. If their vacation lasted for n days, then is the total number of miles they traveled on their vacation. B. The community theater had a capacity of 500 seats. If n is the number of tickets sold for a performance, then is number of seats that were left empty for that performance. C. Jody had 500 dollars in her savings account. She withdrew n dollars to buy some camping supplies. She then had dollars left in her savings account. D. A store had 500 apples to be placed in bags with an equal number of apples in each bag. If n is the number of bags to be used, then is the number of apples in each bag

OpenStudy (anonymous):

500/n

OpenStudy (dan815):

A)=500*n

OpenStudy (anonymous):

thank u

OpenStudy (dan815):

B)=500-n

OpenStudy (dan815):

c)=500-n

OpenStudy (dan815):

D)=500/n

OpenStudy (dan815):

D is right

OpenStudy (anonymous):

oh thanks

OpenStudy (dan815):

u r welcome

OpenStudy (anonymous):

before u leave can u help me on another 1?

OpenStudy (anonymous):

Which is an everyday situation for the math expression with the context provided? 12 – t; comparing people's ages A. Holly is 12. Nick is t years older than Holly. Nick's age is represented by 12 – t. B. Holly's age is t. Nick is 12 years older than Holly. Nick's age is represented by 12 – t. C. Holly is 12. Nick is t years younger than Holly. Nick's age is represented by 12 – t. D. Holly's age is t. Nick is 12 years younger than Holly. Nick's age is represented by 12 – t.